Tickets are numbered 1 to 50 and placed in a box. Three tickets are drawn at random without replacement. What is the probability that their numbers are all greater than 25?
Answer by Nate(3500)

There are 24 numbers that are greater than 25 in the box.
(24/50)(23/49)(22/48) = 12144/117600 = 1012/9800 = 506/4900 = 253/2450
About 10.33%